# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ XI Conclusion
In this paper, we conducted a comprehensive literature review on the application of LLM and LLM-based agents in software engineering. We categorized software engineering into six topics: requirement engineering and documentation, code generation and software development, autonomous learning and decision making, software design and evaluation, software test generation, and software security and maintenance. For each topic, we analyzed the tasks, benchmarks, and evaluation metrics, distinguishing between LLM and LLM-based agents and discussing the differences and impacts they bring. We further analyzed and discussed the models used in the experiments of the 117 collected papers. Additionally, we provided statistics and distinctions between LLM and LLM-based agents regarding datasets and evaluation metrics. The analysis revealed that the emergence of LLM-based agents has led to extensive research and applications across various software engineering topics, demonstrating different emphases compared to traditional LLMs in terms of tasks, benchmarks, and evaluation metrics.

### v2

@@ XI Conclusion
In this paper, we conducted a comprehensive literature review on the application of LLM and LLM-based agents in software engineering. We categorized software engineering into six topics: requirement engineering and documentation, code generation and software development, autonomous learning and decision making, software design and evaluation, software test generation, and software security and maintenance. For each topic, we analyzed the tasks, benchmarks, and evaluation metrics, distinguishing between LLM and LLM-based agents and discussing the differences and impacts they bring. We further analyzed and discussed the models used in the experiments of the 139 collected papers. Additionally, we provided statistics and distinctions between LLM and LLM-based agents regarding datasets and evaluation metrics. The analysis revealed that the emergence of LLM-based agents has led to extensive research and applications across various software engineering topics, demonstrating different emphases compared to traditional LLMs in terms of tasks, benchmarks, and evaluation metrics.